# Orbital Stability of Closely-Spaced Four-planet Systems

### v1

@@ 1 Introduction
Over the past two decades, $\sim 1000$ multi-planet systems have been discovered, most by NASA’s Kepler and TESS spacecraft and radial velocity observations. Analysis of photometric data from Kepler’s prime mission has revealed 709 candidate multi-planet systems, many of which include planets on closely-spaced orbits [10]. Pioneering numerical calculations of the stability of multi-planet systems were done at the dawn of the exoplanet era by [3], followed by numerous subsequent studies, including several that focused on systems of three or five Earth-mass planets orbiting a solar mass star as in [17], [11], and [1]. [18] studied systems with many co-orbital planets. However, the stability of four-planet systems has not been studied extensively. We numerically investigate the orbital dynamics of tens of thousands of hypothetical four-planet systems consisting of Earth-mass planets on initially-circular, geometrically-spaced, coplanar orbits around a star of one solar mass over a wide range of initial semimajor axes spacings and with several different initial longitudinal prescriptions. This is done to also study the effects of initial longitudes on the system lifetimes and thereby build upon the investigation of [9].
@@ 2.1 Initial Semimajor Axes
$a_{j}+\beta\left(\frac{M\textsubscript{$\bigoplus$}}{12(M\textsubscript{$\bigodot$})+(j-1)M\textsubscript{$\bigoplus$})}\right)^{\frac{1}{3}}\frac{(a_{j}+a_{j+1})}{2}=a_{j+1}.$ (5)
@@ 2.3 Integrator and End Conditions
Since two planets passing within 0.01 AU, which is approximately the mutual Hill sphere of two planet each of one Earth mass, of each other would cause a large orbital perturbation and rapid subsequent orbital stability degradation, a close encounter condition was selected as one termination condition. Furthermore, there is also the possibility that the planets would collide, and the WHFast integrator neglects the full effects of very close bodies due to the underlying assumptions of the model as investigated by [19].
@@
Ultimately, the close encounter condition was selected owing to improved simulation speed with nearly identical results. A drawback of this method is that it can fail to detect when the outer planet becomes ejected from the system with minimal perturbations to the other planets. While rare, this anomaly occurred at a mutual Hill radius of 3.493 in the four-planet SL09 dataset and caused a spuriously long-lived system, due to the residual three-planet system being considerably more stable than its four-planet counterpart as will be discussed in Section 2.4. To combat this, a planetary escape boundary was applied at a distance of double of the initial semimajor axis of the planet farthest from the star. However, the vast majority of systems did not terminate due to planetary ejection.
@@ 2.4 Shifted Linear Regression
$\log t_{\rm c}=b\beta+c.$ (7)
@@
$\log t_{\text{c}}=b^{\prime}(\beta-2\sqrt{3})+c^{\prime}.$ (8)
@@ 3.1 SL09 Longitudes
Despite the limited resolution in the range [8.5, 10], we observe the influence of a broad and significant reduction in system lifetimes near the $7/6$ MMR between neighboring planets. System lifetimes near this local minimum are at least four orders of magnitude less than most similarly-spaced systems. Additional resonance-induced reductions between neighboring planets are observed for the second-order $13/11$ MMR and third-order $17/14$ and $19/16$ MMRs. The lifetime reduction stemming from a second-order $7/5$ MMR between non-neighboring planets with one planet orbiting between resonating pairs is also evident. We find that both the affected range of planetary spacings and the extent of lifetime reduction diminish for higher order MMRs and MMRs between non-neighboring planets, as discussed further in Section 3.6. We do not see evidence for the second-order $5/3$ MMR between inner and outer planets, though this may be due to the combined expected weakness of this resonance and the fact that our integrations were halted at $10^{10}$ years.
Figure 2: The top panel shows simulated lifetimes of four-planet systems with SL09 (grey) and Aligned (blue) initial longitudes with MMRs overlain. The bottom panel displays the simulated lifetimes for the Random (purple), $\Delta_{10}$ (red), and Hexagonal (light blue) systems. Here and elsewhere in this work, the order of the overlaid plots corresponds to the order in the legend, as can be seen here with the grey SL09 points on top of the blue Aligned points. Period ratios between neighboring planets are marked above each panel. All results are plotted with a resolution of 0.001 $R_{H}$ (mutual Hill radii, see Equation 4) for systems with $\beta\leq 8.5$, above which system lifetimes are plotted at a resolution of 0.01 up to $\beta=10$ for SL09 longitudes and $\beta=8.7$ for Aligned longitudes. Likewise, the resolution was 0.001 $R_{H}$ for Random longitudes up to $\beta=8.5$, while $\Delta_{10}$ and Hexagonal were completed up to $\beta=8.2$. Upward pointing triangles represent systems which remained stable at $10^{10}$ years, when the integrations were terminated. The vertical lines correspond to the regions of first-order (solid lines), second-order (dashed lines), and third-order (dotted lines) MMRs of neighboring planets (blue), pairs with a single planet orbiting between (purple), and the innermost planet with the outermost planet (black). Note that some resonances near the two-planet stability limit do not have callout boxes. No clear pattern of reduction was observed in the lifetimes of systems near second-order resonances between the innermost and outermost planets nor third-order resonances between neighboring planets for systems more tightly-spaced than $\beta\approx 8.52$, so these resonances are not displayed for small values of $\beta$. Third-order resonances between neighbors are labeled beginning at $26/23$, with the first clear effect observed at $20/17$.
@@ 4 Comparison with Lifetimes of Three-planet and Five-planet Systems
We compare our numerical results against the theoretical results of [12] in Figure 5. For the three-planet case, it was shown that, when considering MMRs, the problem can be reduced to unidimensional Chirikov diffusion. However, this is not extendable to systems with greater number of planets except by using a scaling factor to approximate the effects of the other planets. We use a factor of $K=1.5$, midway between $K=1$ for three-planet systems and $K=2$ for five-planet systems, to represent the four-planet system. We can further numerically confirm that the theoretical results from [12] can model the system lifetimes of four- and five-planet systems. Notably, however, there is a stark inconsistency near a period ratio of 1.14 for the three-planet case that continues for the four- and five-planet cases. This is due to a failure of the MMR overlap criterion as mentioned in [12]. Nonetheless, we see further confirmation that the lifetimes of four-planet systems are closer to those of five-planet systems than three-planet systems, as previously discussed.
@@ 5 Phase-Shifting for Differing Initial Longitudes
An unexpected result of our investigations into four-planet systems is the appearance of a “phase shift” displacement of lifetime extrema towards less widely-spaced initial separations for systems with initially aligned longitudes relative to those with Random (as well as most other) initial longitudes. [8] looked into the instability time distributions for various randomly drawn systems. However, we seek to investigate if there is an underlying mechanism that influences the results of the aforementioned study. The lifetime displacement between longitude sets is evident by comparing the locations of the resonance-induced dips as seen in Figure 2. To investigate the apparent phase-shifting, we use an extrema discovery method detailed in B.1. This method is applied to the relationship between log lifetimes and initial orbital separation for four-planet systems in Section 5.1, allowing us to determine the locations of the various local extrema labeled in Figure 6. It is visually evident that systems with Aligned initial longitudes have lifetime extrema shifted inward (to smaller $\beta$) compared with lifetime extrema locations for well-separated systems with SL09 longitudes. Additionally, this shifting appears to decrease in magnitude as the planetary spacing ($\beta$) increases. Upon comparison with other initial longitude prescriptions, a general progressive trend in phase-shifting is observed as initial longitudes approach conjunction. The physical mechanism responsible for the shift is explained in Section 5.2.
@@ 5.2 Mechanism Responsible for Phase-Shifting
where the first subscript denotes the planet. The second subscript of 0 denotes the initial condition, whereas that of 1000 indicates the average orbital period of the planet measured over the first 1000 years of evolution. Using this measure, we aggregate the change in the periods of the planets from the initial conditions until a preset integration time is reached. The change in the period ratio between four-planet systems with well-spaced and closely-spaced initial longitudes can be seen in Figure 7. Note that the closely-spaced Aligned longitudes are consistently shifted to a higher period ratio, but the shift diminishes as the initial separation grows. For both SL09 and Aligned initial longitudes, there are significant increases in the period ratio that correspond to first-order resonances, especially between planets that are more closely-spaced. Additionally, small positive period shifts coincide with second-order MMRs between neighboring planets for the SL09 initial longitudes.
@@ B.3 LOWESS Smoothing
Robust Locally Weighted Regression and Smoothing Scatterplots (LOWESS), developed by [4], is a robust optimal polynomial fitting technique for smoothing noisy data. This technique is useful because it results in a differentiable relationship from heavily scattered data while maintaining the local behavior. This is accomplished by weighting an aggregation of local polynomial fits in the region. There are also robustness constraints to reject outliers that may occur in order to better represent the bulk trend.

### v2

@@ 1 Introduction
Over the past two decades, $\sim 1000$ multi-planet systems have been discovered, most by NASA’s Kepler and TESS spacecraft and radial velocity observations. Analysis of photometric data from Kepler’s prime mission has revealed 709 candidate multi-planet systems, many of which include planets on closely-spaced orbits [10]. Why multi-planetary systems of this nature form has been explored through planet-planet scattering models by [19] and [17]. The stability impacts of planet-planet scattering are investigated by [12]. In [18] in situ accretion, type 1 migration, gas giant shepherding, and secular resonance shepherding formation models were investigated for close-in planets. Moreover, an analysis of the formation of compact giant planets was completed by [11]. Pioneering numerical calculations of the stability of multi-planet systems were done at the dawn of the exoplanet era by [3], followed by numerous subsequent studies, including several that focused on systems of three or five Earth-mass planets orbiting a solar mass star, such as [24], [13], and [1]. [25] studied systems with many co-orbital planets. However, the stability of four-planet systems has not been studied extensively, although [23] considered four-planet systems with non-equal masses. We numerically investigate the orbital dynamics of tens of thousands of hypothetical four-planet systems consisting of Earth-mass planets on initially-circular, geometrically-spaced, coplanar orbits around a star of one solar mass over a wide range of initial semimajor axes spacings and with several different initial longitudinal prescriptions. This is done to also study the effects of initial longitudes on the system lifetimes and thereby build upon the investigation of [9].
@@ 2.1 Initial Semimajor Axes
$a_{j}+\beta\left(\frac{M\textsubscript{$\bigoplus$}}{12(M\textsubscript{$\bigodot$}+(j-1)M\textsubscript{$\bigoplus$})}\right)^{\frac{1}{3}}\frac{(a_{j}+a_{j+1})}{2}=a_{j+1}.$ (5)
@@ 2.3 Integrator and End Conditions
Since two planets passing within 0.01 AU, which is approximately the mutual Hill sphere of two planets each of one Earth mass, of each other would cause a large orbital perturbation and rapid subsequent orbital stability degradation, a close encounter condition was selected as one termination condition. The presence of destabilizing close encounters are expected after formation as indicated by [4]. Furthermore, there is also the possibility that the planets would collide, and the WHFast integrator neglects the full effects of very close bodies due to the underlying assumptions of the model as investigated by [26].
@@
Ultimately, the close encounter condition was selected owing to improved simulation speed with nearly identical results. A drawback of this method is that it can fail to detect when the outer planet becomes ejected from the system with minimal perturbations to the other planets. While rare, this anomaly occurred at a mutual Hill radius of 3.493 in the four-planet SL09 dataset and caused a spuriously long-lived system, due to the residual three-planet system being considerably more stable than its four-planet counterpart as will be discussed in Section 2.4. To combat this, a planetary escape boundary was applied at a distance of double of the initial semimajor axis of the planet farthest from the star. However, the vast majority of systems did not terminate due to planetary ejection. Planetary ejections of outer planets has been observed in [16].
@@ 2.4 Shifted Linear Regression
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log t_{\rm c}=b\beta+c.$} (7)
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log t_{\text{c}}=b^{\prime}(\beta-2\sqrt{3})+c^{\prime}.$} (8)
@@ 3.1 SL09 Longitudes
Despite the limited resolution in the range [8.5, 10], we observe the influence of a broad and significant reduction in system lifetimes near the $7/6$ MMR between neighboring planets. System lifetimes near this local minimum are at least four orders of magnitude less than most similarly-spaced systems. Additional resonance-induced reductions between neighboring planets are observed for the second-order $13/11$ MMR and third-order $19/16$ and $20/17$ MMRs. The lifetime reduction stemming from a second-order $7/5$ MMR between non-neighboring planets with one planet orbiting between resonating pairs is also evident. We find that both the affected range of planetary spacings and the extent of lifetime reduction diminish for higher order MMRs and MMRs between non-neighboring planets, as discussed further in Section 3.6. We do not see evidence for the second-order $5/3$ MMR between inner and outer planets, though this may be due to the combined expected weakness of this resonance and the fact that our integrations were halted at $10^{10}$ years.
Figure 2: The top panel shows the lifetimes of four-planet systems with SL09 (grey) and Aligned (blue) initial longitudes with MMRs overlain. The bottom panel displays the simulated lifetimes for the Random (purple), $\Delta_{10}$ (red), and Hexagonal (light blue) systems. Here and elsewhere in this work, the order of the overlaid plots corresponds to the order in the legend, as can be seen here with the grey SL09 points on top of the blue Aligned points. Period ratios between neighboring planets are marked above each panel. All results are plotted with a resolution of 0.001 $R_{H}$ (mutual Hill radii, see Equation 4) for systems with $\beta\leq 8.5$, above which system lifetimes are plotted at a resolution of 0.01 up to $\beta=10$ for SL09 longitudes and $\beta=8.7$ for Aligned longitudes. Likewise, the resolution was 0.001 $R_{H}$ for Random longitudes up to $\beta=8.5$, while $\Delta_{10}$ and Hexagonal were completed up to $\beta=8.2$. Upward pointing triangles represent systems which remained stable at $10^{10}$ years, when the integrations were terminated. The vertical lines correspond to the regions of first-order (solid lines), second-order (dashed lines), and third-order (dotted lines) MMRs of neighboring planets (blue), pairs with a single planet orbiting between (purple), and the innermost planet with the outermost planet (black). Note that some resonances near the two-planet stability limit do not have callout boxes. No clear pattern of reduction was observed in the lifetimes of systems near second-order resonances between the innermost and outermost planets nor third-order resonances between neighboring planets for systems more tightly-spaced than $\beta\approx 8.52$, so these resonances are not displayed for small values of $\beta$. Third-order resonances between neighbors are labeled beginning at $26/23$, with the first clear effect observed at $20/17$.
@@ 4 Comparison with Lifetimes of Three-planet and Five-planet Systems
We compare our numerical results against the theoretical results of [14] in Figure 5. For the three-planet case, it was shown that, when considering MMRs, the problem can be reduced to unidimensional Chirikov diffusion. However, this is not extendable to systems with greater number of planets except by using a scaling factor to approximate the effects of the other planets. We use a factor of $K=1.5$, midway between $K=1$ for three-planet systems and $K=2$ for five-planet systems in [14], to represent the four-planet system. We can further numerically confirm that the theoretical results from [14] can model the system lifetimes of four- and five-planet systems. Notably, however, there is a stark inconsistency near a period ratio of 1.14 for the three-planet case that continues for the four- and five-planet cases. This is due to a failure of the MMR overlap criterion as mentioned in [14]. Nonetheless, we see further confirmation that the lifetimes of four-planet systems are closer to those of five-planet systems than three-planet systems, as previously discussed.
@@ 5 Phase-Shifting for Differing Initial Longitudes
An unexpected result of our investigations into four-planet systems is the appearance of a “phase shift” displacement of lifetime extrema towards less widely-spaced initial separations for systems with initially Aligned longitudes relative to those with Random (as well as most other) initial longitudes. The lifetime displacement between longitude sets is evident by comparing the locations of the resonance-induced dips as seen in Figure 2. To investigate the apparent phase-shifting, we use an extrema discovery method detailed in B.1. This method is applied to the relationship between log lifetimes and initial orbital separation for four-planet systems in Section 5.1, allowing us to determine the locations of the various local extrema labeled in Figure 6. It is visually evident that systems with Aligned initial longitudes have lifetime extrema shifted inward (to smaller $\beta$) compared with lifetime extrema locations for well-separated systems with SL09 longitudes. Additionally, this shifting appears to decrease in magnitude as the planetary spacing ($\beta$) increases. Upon comparison with other initial longitude prescriptions, a general progressive trend in phase-shifting is observed as initial longitudes approach conjunction. The physical mechanism responsible for the shift is explained in Section 5.2.
@@ 5.2 Mechanism Responsible for Phase-Shifting
where the first subscript denotes the planet. The second subscript of 0 denotes the initial condition, whereas that of 1000 indicates the average orbital period of the planet measured over the first 1000 years of evolution. Using this measure, we aggregate the change in the periods of the planets from the initial conditions until a preset integration time is reached. The change in the period ratio between four-planet systems with well-spaced and closely-spaced initial longitudes can be seen in Figure 7. Note that the closely-spaced Aligned longitudes are consistently shifted to a higher period ratio, but the shift diminishes as the initial separation grows. This trend corresponds closely with the results in Table 3 wherein the difference in period ratio for the phase shifting. This is exemplified by considering an initial separation of approximately 5.5, $5,\,7.0,\,\text{and }8.0$ mutual Hill radii with a period shifting of $0.775\times 10^{-3},\,0.548\times 10^{-3},\,\text{and }0.284\times 10^{-3}$. For both SL09 and Aligned initial longitudes, there are significant increases in the period ratio that correspond to first-order resonances, especially between planets that are more closely-spaced. Additionally, small positive period shifts coincide with second-order MMRs between neighboring planets for the SL09 initial longitudes.
@@ B.3 LOWESS Smoothing
Robust Locally Weighted Regression and Smoothing Scatterplots (LOWESS), developed by [5], is a robust optimal polynomial fitting technique for smoothing noisy data. This technique is useful because it results in a differentiable relationship from heavily scattered data while maintaining the local behavior. This is accomplished by weighting an aggregation of local polynomial fits in the region. There are also robustness constraints to reject outliers that may occur, in order to better represent the bulk trend.